OpenStudy (anonymous):

Round the answer to the nearest cent. 5% 1/2 tax on $20 is $____.a0 please help me:(

jimthompson5910 (jim_thompson5910):

5 and 1/2 = 5 + 1/2 5 and 1/2 = 5 + 0.5 5 and 1/2 = 5.5 So 5 and 1/2% = 5.5% Divide by 100 to convert to decimal form 5.5% ---> 5.5/100 = 0.055 So 5 and a half percent converts to 0.055 Once you have this decimal value, you multiply it by 20 to get the amount you have to pay in taxes You then add this amount in taxes to the original price.

jimthompson5910 (jim_thompson5910):

oh wait, disregard the last line, they just want the taxes only I think